OpenStudy (anonymous):

Algebra 2 help with focal length

OpenStudy (amistre64):

you guessed eh ....

OpenStudy (anonymous):

yeah i know the steps but not sure how to solve the steps

OpenStudy (amistre64):

we have a focus and a directrix, what do these terms mean to you?

OpenStudy (anonymous):

focus is a point and the directrix is the line across from it i think thats how you describe it

OpenStudy (amistre64):



OpenStudy (amistre64):

since we can use a vertex to define the situation, the vertex is the midpoint between the focus and directrix

OpenStudy (anonymous):

Thats where i messed up i dont know how to find the midpoint

OpenStudy (amistre64):

the points we are looking for all share the same x component in this case soo f(-5, 5) <------ v(-5,(5-1)/2) d(-5,-1)

OpenStudy (amistre64):

so we have some form of y = a(x-h)^2 + k y = a(x+5)^2 + 2

OpenStudy (amistre64):

since the focus is above the directrix, we are open up so a -a is not going to fit ...

OpenStudy (anonymous):

wait when you said d(-5,-1) was that the directrix?

OpenStudy (amistre64):

that was the point in the directrix that is in line with the focus and vertex

OpenStudy (anonymous):

so thats the bottom point in the drawing

OpenStudy (amistre64):

the directrix is a line thru y-1; so all points on the directrix are of the form (x,-1) and since all the points in line fvd have the same x value ... d(-5,-1) is it yes, thats the point in the picture

OpenStudy (amistre64):

**y=-1

OpenStudy (anonymous):

what was the vertex again?